# Generating variable $\hbar$ and $c$ via Fujii–Wetterich model in curved spacetimes

### v1

@@ II Phenomenology of VSL
in which the speed of light $c$ is a function of the scalon field $\chi$ (per $c\propto\chi^{1/2}$ as specified in Eq. (21)). In the pure $\mathcal{R}^{2}$ gravity theory, Eq. (12), the scalon field is directly related to the Ricci scalar as $\mathcal{R}=2f_{0}\,\chi^{2}$, whereas for the metric (50) it can be shown that $\mathcal{R}\propto a^{-2}$. Consequently, in the cosmic void where the generalized FLRW metric is applicable, the following relation holds  [^12] This relation is a special case of a more generic form $c\propto a^{-\zeta}$ which Barrow heuristically proposed in [40]. In [15] we consider Barrow’s form in the context of power-law cosmology [41, 42].
@@
$ds^{2}=\frac{c_{0}^{2}}{a(t)}\,dt^{2}-a^{2}(t)\,\left(\frac{dr^{2}}{1-\kappa\,r^{2}}+r^{2}\,d\Omega^{2}\right)$ (53)
where the cosmic scale factor at our current time $t_{0}$ is set equal to 1, and $c_{0}$ is the speed of light measured in the void at $t_{0}$. It is important to note that several authors have previously applied a VSL cosmology to SnIa standard candles, while circumventing dark energy [43, 46, 44, 49, 48, 50, 47, 45, 51]. The consensus among these studies is that VSL alone could not adequately account for the SnIa data. However, these conclusions warrant reconsideration, as all previous analyses implicitly assumed the speed of light depends solely on cosmic time $t$. This assumption fails to hold for our VSL cosmology whereby $c$ varies in both space and time, making it location-dependent. This is because the scalon field $\chi$ can vary in spacetime, with its value in the intergalactic void (which is subject to cosmic expansion) differing from that in gravitationally bound galaxies (which host SnIa and are resistant to cosmic expansion). Hence, as $c\propto\chi^{1/2}$, $c$ is not only time-dependent but also location-dependent. A thorough treatment of this intricate issue is provided in our companion paper [14].
In the presence of VSL, the classic Lemaître redshift formula, $1+z=a^{-1}$, in standard cosmology is no longer applicable. Previous analyses of SnIa using VSL overlooked this crucial distinction and continued to use the classic formula, resulting in incorrect conclusions. The dependence of $c$ on the scalon field $\chi$ introduces two fundamental modifications to the right hand side of the Lemaître redshift relation, as follows:
@@
    An additional correction accounts for the variation of the scalon field among galaxies as the function of redshift $z$. This is because galaxies hosting SnIa may possess different values for the scalon field due to their evolution in the expanding cosmic background. We model this physically motivated effect by introducing a monotonic function $F(z)$ that smoothly interpolates between two extremes: $F(z\rightarrow 0)=1$ and $F(z\rightarrow\infty)=F_{\infty}$ (with $F_{\infty}$ being an adjustable parameter). Specifically, we find the following formulation to be suitable
@@
    By virtue of the length scaling (26), $F(z)$ thus signifies the evolution in the typical size of galaxies in response to the expansion of the comic void. This astronomical effect will have implications for the $H_{0}$ tension, which will be shown later in Point VIII.
Derivations of these modifications are detailed in our companion paper [14]. Here, we only quote the results. In our VSL cosmology, the modified Lemaître redshift formula reads
@@
This significantly contrasts with the usual Hubble law $z\approx H_{0}\,d/c$ in standard cosmology. The multiplicative $3/2$–factor in Eq. (56) will have crucial implications in the estimation of $H_{0}$, as we shall see in Points III and V.
In a scale-invariant cosmology, the cosmic scale factor evolves as
$a=\left(\frac{t}{t_{0}}\right)^{2/3}$ (57)
reflecting the generic time scaling rule, $\tau\propto l^{3/2}$, derived in Eq. (27). Note that this evolution also occurs in the flat $\Lambda$CDM model in absence of dark energy ($\Omega_{\Lambda}=0$). Combining (55) and (57), we obtain the modified luminosity distance–redshift relation applicable for our VSL cosmology [14]
@@
For comparison, this formula fundamentally differs from the canonical ones in the flat $\Lambda$CDM model ($\Omega_{M}+\Omega_{\Lambda}\!=\!1$)
@@
and in the flat Einstein–de Sitter (EdS) universe (i.e. $\Omega_{M}=1,\,\Omega_{\Lambda}=0$)
$\frac{d_{L}}{c}=2\,\frac{1+z}{H_{0}}\left(1-\frac{1}{\sqrt{1+z}}\right)\,.$ (60)
II. Analyzing Pantheon Catalog of SnIa using VSL—The modified $d_{L}$-vs-$z$ relation (58), supplemented with (54), now stands ready for application to the Pantheon Catalog of SnIa. Our detailed analysis is provided in Ref. [14]. The Pantheon dataset, as provided in [52], is shown as open circles in the Hubble diagram in Fig. 1. It comprises $N=1,048$ data points given in terms of redshift and luminosity distance $\{z,\,d_{L}\}$, together with an error bar $\sigma_{L}$ of $d_{L}$, for $z$ spanning the range (0, 2.26)02.26(0,\,2.26)( 0 , 2.26 ). Fora given theoretical model, one would minimize the error
$\chi^{2}:=\frac{1}{N}\sum_{i=1}^{N}\biggl{(}\frac{d_{L}^{\text{model}}-d_{L}^{\text{Pantheon}}}{\sigma_{L}^{\text{Pantheon}}}\biggr{)}^{2}\,.$ (61)
Figure 1: Hubble diagram of SnIa in the Pantheon Catalog. Open circles: 1,048 data points with error bars, listed in Ref. [52]. In both panels, long-dashed line is $\Lambda$CDM Formula (59) with $H_{0}=70.2$, $\Omega_{M}=0.285$, $\Omega_{\Lambda}=0.715$; dotted line is flat CDM Formula (60) with $H_{0}=70.2$. Upper panel: Solid line is our VSL Formula (58) with $H_{0}=47.2$ and $F_{\infty}=0.93$. Lower panel: Solid line is our VSL Formula (58) with $F(z)\equiv 1\ \forall z$ with $H_{0}=44.4$.
Let us first review the canonical interpretation of SnIa data in standard cosmology. For very low $z$, both formulae (59) and (60) yield $\frac{d_{L}}{c}\approx\frac{z}{H_{0}}$. However, for positive $\Omega_{\Lambda}$, given the same $H_{0}$, the flat $\Lambda$CDM formula (59) produces a higher value for $d_{L}$ compared to the flat EdS formula (60). The crossover occurs at a redshift $z_{*}$ which approximately satisfies $\Omega_{M}(1+z_{*})^{3}\approx\Omega_{\Lambda}$. We fit Formula (59) to the Pantheon Catalog and obtain $H_{0}=70.2$, $\Omega_{M}=0.285$, $\Omega_{\Lambda}=0.715$ with an error $\chi_{\text{min}}^{2}=0.98824$. Hence, starting at $z\gtrsim z_{*}\approx 0.36$, the flat $\Lambda$CDM model begins to show an excess in $d_{L}$, meaning high-$z$ SnIa’s appear dimmer than the flat EdS model would predict. This behavior, manifest in the high-$z$ portion of the Hubble diagram (Fig. 1), has been interpreted as evidence for late-time cosmic accelerating expansion [58, 59] and for the existence of dark energy.
@@
This indicates that our analysis practically—and parsimoniously—requires only one parameter $H_{0}$. The function $F(z)$ plays a role in resolving the $H_{0}$ tension, which will be explained in Point VIII below.
[…]
One pertinent question is the origin of this reduction in $H_{0}$. As we explained in Point I, for very low $z$, the Hubble law for our VSL cosmology is modified to $z\approx\frac{3}{2}H_{0}\,d/c_{0}$, as opposed to the standard Hubble law $z\approx H_{0}\,d/c$. Thus, for our VSL cosmology, the numerical value (empirically known to be $\sim 71$) for the slope of the $z$-vs-$d$ relation is not $H_{0}$ but $\frac{3}{2}H_{0}$, which promptly leads to $H_{0}\sim 47$. In conclusion, it is the $3/2-$exponent in the $a^{-3/2}$ term in our modified Lemaître redshift formula (55) that is responsible for the reduction in the $H_{0}$ value by a factor of $3/2$ compared to the canonical range.
Surprisingly, this reduced value of $H_{0}$ finds corroboration in a remarkable discover made by Blanchard et al in 2023 [54], when these authors analyzed the CMB power spectrum. We shall review their work below.
IV. Revitalizing Blanchard et al (BDRS)’s analysis of CMB power spectrum, bypassing DE—It is well established that the $\Lambda$CDM, with the primordial fluctuations spectrum $P(k)\propto k^{n}$, successfully accounts for Planck’s thermal power spectrum of the Cosmic Microwave Background (CMB) with parameters $n\approx 0.96$, $\Omega_{\Lambda}\approx 0.7$, $H_{0}\approx 67$. However, in 2003, Blanchard, Douspis, Rowan-Robinson, and Sarkar (BDRS) made a surprising discovery [54]. Instead of adhering to the standard spectrum $P(k)\propto k^{n}$, they considered two separate power-law forms for the low $k$ and high $k$ regimes, viz.
@@
where $k_{*}$ is a break point, and continuity across $k_{*}$ is enforced by the condition $A_{1}\,k_{*}^{n_{1}}=A_{2}\,k_{*}^{n_{2}}$. Most importantly, they deliberately set $\Omega_{\Lambda}=0$, thereby avoiding the dark energy hypothesis. Remarkably, their model achieved an excellent fit to WMAP’s CMB data available at the time, with the optimal parameter values including $H_{0}=46$, $k_{*}=0.0096$ Mpc-1, $n_{1}=1.015$, $n_{2}=0.806$ [54]. Of these parameters, the most striking aspect was the low Hubble value $H_{0}=46$. A subsequent in-depth study by Hunt and Sarkar [55], utilizing a supergravity-based multiple inflation scenario, also successfully accounted for the CMB power spectrum without invoking dark energy, while requiring a similar value of $H_{0}\approx 43.5$. (Interestingly, Shanks in 2004 [56] also ventured that if $H_{0}\lesssim 50$ then a simpler, inflationary model with $\Omega_{baryon}=1$ might be allowed with no need for dark energy or cold dark matter.)
@@
It is also important to note that the perceived low $H_{0}$ value $\sim 46$ eventually led BDRS to abandon their original findings in a follow-up study related to SDSS’s two-point correlation data of luminous red galaxies [57]. However, their abandonment may have been premature, since in the VSL framework the anomalous $3/2$–exponent in the anisotropic time scaling influences the observation of distant light sources along the time direction. Therefore, a reassessment of BDRS’s work of SDSS data, incorporating this effect, would be warranted. Additionally, future investigations should explore the impacts of VSL on phenomena such as lensing, Baryon Acoustic Oscillations (BAO), and the Etherington distance duality relation.
V. Resolving the age problem, bypassing DE—The flat $\Lambda$CDM model gives the age formula
@@
VI. Obtaining late-time acceleration via VSL, bypassing DEIt is an empirical fact that high-$z$ SnIa’s appear fainter than what would be predicted by the (matter-dominated) flat EdS model [58, 59]. This phenomenon is reflected in the distance modulus excess observed in the high-$z$ segment of the Hubble diagram in Fig. 1. Standard cosmology attributes this behavior to late-time cosmic acceleration, necessitating the introduction of a cosmological constant $\Lambda$ into the Friedmann equations, resulting in the $\Lambda$CDM model where $\Omega_{\Lambda}$ quantifies the amount of dark energy.
However, the analysis presented in Points I and II above demonstrates the efficacy of our VSL cosmology in accounting for the excess in $d_{L}$ without invoking dark energy. Its success can be quantitatively understood as follows. At high $z$, the luminosity distance formula of the flat EdS model (60) behaves as

### v2

@@ VI.1 VSL cosmology and late-time cosmography
by allowing the speed of light $c$ to vary alongside the dilaton field $\chi$ (per $c\propto\chi^{1/2}$ as specified in Eq. (21)). Since the dilaton directly determines the lengthscale for physical processes, $\chi\propto l^{-1}$, as in Eq. (26), and the cosmic scale factor plays the role of a lengthscale in the FLRW metric, it is reasonable to make the following ansatz:
Ansatz #1:
$\chi\propto a^{-1}$ (51)
Furthermore, since the dilaton also determines the timescale for physical processes, $\chi\propto\tau^{-2/3}$, as in Eq. (27), and the cosmic time $t$ plays the role of a timescale in the FLRW metric, it is reasonable to make the following ansatz on the evolution of the cosmic scale factor:
Ansatz #2:
[…]
This ansatz is identical to the evolution of the scale factor in the Einstein–de Sitter (EdS) universe. Obviously, the vanilla EdS model—by itself—cannot account for the late-time cosmic acceleration observed in the Hubble diagram of Type Ia supernova (SNeIa). This failure forms the basis to replace the EdS model by the “concordance” cosmological model in which a $\Lambda$ component of density $\Omega_{\Lambda}\approx 0.7$.
However, in the rest of this paper, we consider the EdS universe in conjunction with a varying speed of light. From $c\propto\chi^{1/2}$ and Ansatz #1, we deduce that, in the intergalactic space, the following relation holds
@@
$ds^{2}=\frac{c_{0}^{2}}{a(t)}\,dt^{2}-a^{2}(t)\,\biggl{(}\frac{dr^{2}}{1-\kappa\,r^{2}}+r^{2}\,d\Omega^{2}\biggr{)}$ (54)
where the cosmic scale factor at our current time $t_{0}$ is set equal to 1 (i.e., $a_{0}=1$), and $c_{0}$ is the speed of light measured in the intergalactic space at $t_{0}$. It is important to note that several authors have previously applied a VSL cosmology to SNeIa standard candles, while circumventing dark energy [26, 27, 32, 29, 28, 33, 31, 30, 34]. The consensus among these studies is that VSL alone could not adequately account for the SNeIa data. However, these conclusions warrant reconsideration, as all previous analyses implicitly assumed the speed of light depends solely on cosmic time $t$. This assumption fails to hold for our VSL cosmology whereby $c$ varies in both space and time, making it location-dependent. This is because the dilaton field $\chi$ can vary in spacetime, with its value in the intergalactic space (which is subject to cosmic expansion) differing from that in gravitationally bound galaxies (which host SNeIa and are resistant to cosmic expansion). Hence, as $c\propto\chi^{1/2}$, $c$ is not only time-dependent but also location-dependent. A thorough treatment of this intricate issue is provided in Ref. [19].
In the presence of VSL, the classic Lemaître redshift formula, $1+z=a^{-1}$, in standard cosmology is no longer applicable. Previous analyses of SNeIa using VSL overlooked this crucial distinction and continued to use the classic formula, resulting in incorrect conclusions. The dependence of $c$ on the dilaton field $\chi$ introduces two fundamental modifications to the right hand side of the Lemaître redshift relation, as follows:
@@
    An additional correction accounts for the variation of the dilaton field among galaxies as the function of redshift $z$. This is because galaxies hosting SNeIa may possess different values for the dilaton field due to their evolution in the expanding cosmic background. We model this physically motivated effect by introducing a monotonic function $F(z)$ that smoothly interpolates between two extremes: $F(z\rightarrow 0)=1$ and $F(z\rightarrow\infty)=F_{\infty}$ (with $F_{\infty}$ being an adjustable parameter). Specifically, we find the following formulation to be suitable
@@
    By virtue of the length scaling (26), $F(z)$ thus signifies the evolution in the typical size of galaxies in response to the expansion of the intergalactic space. This astronomical effect will have implications for the $H_{0}$ tension, which will be shown later in Section VI.8.
Derivations of these modifications are detailed in Ref. [19]. Here, we only quote the results. In our VSL cosmology, the modified Lemaître redshift formula reads
@@
This result contrasts with the Hubble law $z\approx H_{0}\,d/c$in standard cosmology. The multiplicative $3/2$–factor in Eq. (57) will have crucial implications in the estimation of $H_{0}$, as we shall see in Sections VI.3 and VI.5.
Combining Eq. (56) and Ansatz #2 (52), we obtain the modified luminosity distance–redshift relation applicable for our VSL cosmology [19]
@@
For comparison, this formula fundamentally differs from the canonical ones in the spatially flat $\Lambda$CDM model ($\Omega_{M}+\Omega_{\Lambda}\!=\!1$)
@@
and in the EdS universe (i.e. $\Omega_{M}=1,\,\Omega_{\Lambda}=0$)
$\frac{d_{L}}{c}=2\,\frac{1+z}{H_{0}}\,\biggl{(}1-\frac{1}{\sqrt{1+z}}\biggr{)}\,.$ (60)
VI.2 Analyzing Pantheon Catalog of SNeIa using VSL
The modified $d_{L}$-vs-$z$ relation (58), supplemented with (55), now stands ready for application to the Pantheon Catalog of SNeIa. Our detailed analysis is provided in Ref. [19]. The Pantheon dataset, as provided in Ref. [139], is shown as open circles in the Hubble diagram in Fig. 1. It comprises $N=1,048$ data points given in terms of redshift and luminosity distance $\{z,\,\mu\}$, together with an error bar $\sigma$ of $\mu$, for $z$ spanning the range (0, 2.26)02.26(0,\,2.26)( 0 , 2.26 ). For a given theoretical model, one would minimize the error
$\mu:=m-M=5\log_{10}(d_{L}/\text{Mpc})+25$ (61)
$\chi^{2}:=\frac{1}{N}\sum_{i=1}^{N}\biggl{(}\frac{\mu_{i}^{\text{model}}-\mu_{i}^{\text{Pantheon}}}{\sigma_{i}^{\text{Pantheon}}}\biggr{)}^{2}\,.$ (62)
Let us first review the canonical interpretation of SNeIa data in standard cosmology. For very low $z$, both formulae (59) and (60) yield $\frac{d_{L}}{c}\approx\frac{z}{H_{0}}$. However, for positive $\Omega_{\Lambda}$, given the same $H_{0}$, the flat $\Lambda$CDM formula (59) produces a higher value for $d_{L}$ compared to the EdS formula (60). The crossover occurs at a redshift $z_{*}$ which approximately satisfies $\Omega_{M}(1+z_{*})^{3}\approx\Omega_{\Lambda}$. We fit Formula (59) to the Pantheon Catalog and obtain $H_{0}=70.2$, $\Omega_{M}=0.285$, $\Omega_{\Lambda}=0.715$ with an error $\chi_{\text{min}}^{2}=0.98824$. Hence, starting at $z\gtrsim z_{*}\approx 0.36$, the spatially flat $\Lambda$CDM model begins to show an excess in $d_{L}$, meaning high-$z$ SNeIa appear dimmer than the EdS model would predict. This behavior, manifest in the high-$z$ portion of the Hubble diagram (Fig. 1), has been interpreted as evidence for late-time cosmic accelerating expansion [140, 141] and for the existence of dark energy.
@@ VI.2 Analyzing Pantheon Catalog of SNeIa using VSL
Figure 1: Hubble diagram of SNeIa in the Pantheon Catalog. Open circles: 1,048 data points with error bars, listed in Ref. [139]. In both panels, long-dashed line is the $\Lambda$CDM Formula (59) with $H_{0}=70.2$, $\Omega_{M}=0.285$, $\Omega_{\Lambda}=0.715$; dotted line is the EdS Formula (60) with $H_{0}=70.2$. Upper panel: Solid line is our VSL Formula (58) with $H_{0}=47.2$ and $F_{\infty}=0.93$. Lower panel: Solid line is our VSL Formula (58) with $F(z)\equiv 1\ \forall z$ with $H_{0}=44.4$.
@@
This indicates that our analysis practically—and parsimoniously—requires only one parameter $H_{0}$. The function $F(z)$ plays a role in resolving the $H_{0}$ tension, which will be explained in Section VI.8 below.
VI.3 A reduced value of $H_{0}$ by a factor of $3/2$
A striking result deduced from our SNeIa analysis thus far is a low value $H_{0}=47.2$, which starkly contrasts with the local measurement $H_{0}=73$ widely accepted in the $\Lambda$CDM framework. Note that this low value is inherent to the modified $d_{L}$-vs-$z$ formula (58) regardless of the function $F(z)$. When $F(z)$ is disabled, the formula still produces $H_{0}=44.4$, a value comparable to that obtained when $F(z)$ is enabled. The posterior distribution of $H_{0}$ is shown in the left panel of Fig. 3 and corresponds to $H_{0}=47.2\pm 0.4$ (at 95% confidence level).
We should note that a method of “redshift remapping” was introduced in [143, 144, 145]. Our modified Lemaître redshift formula could be viewed as effectively a “redshift remapping”. Interestingly, in [145], based on this method, a joint analysis of all primary cosmological probes including the local measurement of the Hubble constant, SNeIa, Baryon Acoustic Oscillations (BAO), Planck observations of the CMB power spectrum and cosmic chronometers yields $H_{0}=48\pm 2$, a range in good agreement with our result $H_{0}=47.2\pm 0.4$.
The drastic deviation in the current-time Hubble value between our $H_{0}=47.2$ from the local measurement $H_{0}=73$ can be understood by using the modified Hubble law (57) in VSL cosmology. For very low $z$, the Hubble law for our VSL cosmology is modified to $z\approx\frac{3}{2}H_{0}\,d/c_{0}$, as opposed to the standard Hubble law $z\approx H_{0}\,d/c$. Thus, for our VSL cosmology, the numerical value (empirically known to be $\sim 71$) for the slope of the $z$-vs-$d$ relation is not $H_{0}$ but $\frac{3}{2}H_{0}$, which promptly leads to $H_{0}\sim 47$. In conclusion, it is the $3/2-$exponent in the $a^{-3/2}$ term in our modified Lemaître redshift formula (56) that is responsible for the reduction in the $H_{0}$ value by a factor of $3/2$ compared to the canonical range.
Surprisingly, this reduced value of $H_{0}$ finds corroboration in a remarkable proposal made by Blanchard et al in 2023 [146], when these authors analyzed the CMB power spectrum. We shall review their work below.
VI.4 Revitalizing Blanchard et al (BDRS)’s analysis of CMB power spectrum, bypassing DE
It is well established that the $\Lambda$CDM, with the primordial fluctuations spectrum $P(k)\propto k^{n}$, successfully accounts for Planck’s thermal power spectrum of the Cosmic Microwave Background (CMB) with parameters $n\approx 0.96$, $\Omega_{\Lambda}\approx 0.7$, $H_{0}\approx 67$. However, in 2003, Blanchard, Douspis, Rowan-Robinson, and Sarkar (BDRS) made a surprising discovery [146]. Instead of adhering to the standard primordial fluctuation spectrum $P(k)\propto k^{n}$, they considered two separate power-law forms for the low $k$ and high $k$ regimes, viz.
@@ VI.4 Revitalizing Blanchard et al (BDRS)’s analysis of CMB power spectrum, bypassing DE
where $k_{*}$ is a break point, and continuity across $k_{*}$ is enforced by imposing $A_{1}\,k_{*}^{n_{1}}=A_{2}\,k_{*}^{n_{2}}$. Most importantly, they deliberately set $\Omega_{\Lambda}=0$, thereby avoiding the dark energy hypothesis. Remarkably, their model achieved an excellent fit to WMAP’s CMB data available at the time, with the optimal parameter values including $H_{0}=46$, $k_{*}=0.0096$ Mpc-1, $n_{1}=1.015$, $n_{2}=0.806$ [146]. Of these parameters, the most striking aspect was the low Hubble value $H_{0}=46$. A subsequent in-depth study by Hunt and Sarkar [147], utilizing a supergravity-based multiple inflation scenario, also successfully accounted for the CMB power spectrum without invoking dark energy, while requiring a similar value of $H_{0}\approx 43.5$. (Interestingly, Shanks in 2004 [148] also ventured that if $H_{0}\lesssim 50$ then a simpler, inflationary model with $\Omega_{baryon}=1$ might be allowed with no need for dark energy or cold dark matter.)
@@
It is also important to note that the perceived low $H_{0}$ value $\sim 46$ eventually led BDRS to abandon their original findings in a follow-up study related to SDSS’s two-point correlation data of luminous red galaxies [149]. However, their abandonment may have been premature, since in the VSL framework the anomalous $3/2$–exponent in the anisotropic time scaling influences the observation of distant light sources along the time direction. Therefore, a reassessment of BDRS’s work of SDSS data, incorporating this effect, would be warranted. Additionally, future investigations should explore the impacts of VSL on phenomena such as lensing, BAO, and the Etherington distance duality relation.
As we stated earlier in Section VI.3, using a “redshift remapping” method for all primary cosmological probes, the authors in [145] obtained $H_{0}=48\pm 2$, which is compatible with our value $H_{0}=47.2$ and the value $H_{0}=46$ obtained by BDRS.
VI.5 Resolving the age problem, bypassing DE
The spatially flat $\Lambda$CDM model gives the age formula
@@ VI.5 Resolving the age problem, bypassing DE
VI.6 Obtaining late-time acceleration via VSL, bypassing DE
Figure 4: The physical intuition of late-time acceleration based on VSL as explained in Section VI.7. In the left panel, photons from SNe A and SNe C were emitted at times $t_{A}$ and $t_{C}$ with $t_{C}-t_{0}=2\,(t_{A}-t_{0})$; thus, their redshifts satisfy $z_{C}=2\,z_{A}$. However, for the photons emitted from SNe C, the earlier segment of their trajectory had a higher speed of light than the later segment, allowing them to cover a longer distance. This results in $d_{C}>2\,d_{A}$, leading in an excess distance modulus in the Hubble diagram for high-$z$ SNe, as depicted in the right panel.
It is an empirical fact that high-$z$ SNeIa’s appear fainter than what would be predicted by the EdS model [140, 141]. This phenomenon is reflected in the distance modulus excess observed in the high-$z$ segment of the Hubble diagram in Fig. 1. Standard cosmology attributes this behavior to late-time cosmic acceleration, necessitating the introduction of a cosmological constant $\Lambda$ into the Friedmann equations, resulting in the $\Lambda$CDM model where $\Omega_{\Lambda}$ quantifies the amount of dark energy.
However, the analysis presented in Sections VI.1 and VI.2 demonstrates the efficacy of our VSL cosmology in accounting for the excess in $d_{L}$ without invoking dark energy. Its success can be quantitatively understood as follows. At high $z$, the luminosity distance formula of the EdS model (60) behaves as